OpenStudy (anonymous):

find inverse laplace transform of F(s)= (s^2+2s+5/(s^2(s+1))

OpenStudy (anonymous):

note

OpenStudy (anonymous):

'note' sorry this is my first time using this what do you mean by that?

OpenStudy (anonymous):

Im just bookmarking this question so that I can return to it later. :D Id want to know the answer too

OpenStudy (anonymous):

oh ok sounds good thanks

OpenStudy (lgbasallote):

your question is \[\huge \mathcal L ^{-1} \left \{ s^2 + 2s + \frac 5{s^2(s+1)}\right \}\]

OpenStudy (lgbasallote):

??

OpenStudy (lgbasallote):

@ragman you there?

OpenStudy (anonymous):

hey sorry about the late reply but my question is actually (s^2+22+5) / (s^2(s+1) sorry i forgot the other )

OpenStudy (anonymous):

@lgbasallote

OpenStudy (lgbasallote):

uhh s^2 + 22 + 5??

OpenStudy (anonymous):

dang i keep doing typos 2s not 22 so yeah its (s^2+2s+5)

OpenStudy (lgbasallote):

so.. \[\huge \mathcal L ^{-1} \left \{ \frac{s^2 + 2s + 5}{s^2(s+1)} \right \}\]

OpenStudy (anonymous):

yup perfect

hartnn (hartnn):

can u do partial fractions ??

OpenStudy (anonymous):

@hartnn yeah thats what i was doing but for some reason i could figure out the other two variable maybe im solving it incorrectly but i got -3 for A and 5 for B

OpenStudy (lgbasallote):

where's C?

hartnn (hartnn):

those are correct, so u need only C right ? where u took C above s+1

OpenStudy (lgbasallote):

your P.F. should be \[\huge \frac As + \frac B{s^2} + \frac C{s+1}\]

OpenStudy (lgbasallote):

so you have \[\implies As(s+1) + B(s+1) + Cs^2 = s^2 + 2s + 5\]

OpenStudy (anonymous):

oh ok so thats where i messed up i thought that where C is it should be (Cs+D) so i was trying to solve for both of those variables

hartnn (hartnn):

nopes only C, if it were s^2+1 u would have taken Cx+D

hartnn (hartnn):

so i think u can get C now, right?

OpenStudy (anonymous):

oh i see now and yeah i believe i can get C now trying it right now

OpenStudy (lgbasallote):

i think it will change your A value too

OpenStudy (lgbasallote):

so you best check

hartnn (hartnn):

great tell us what u get C as....we all will verify that. A is correct.

OpenStudy (lgbasallote):

nope nevermind..it stays the same

OpenStudy (anonymous):

ok sounds good got B=5 finding A now

OpenStudy (anonymous):

A=-3 and C=4 is that what you guys got?

hartnn (hartnn):

yup, same.Good work.

OpenStudy (lgbasallote):

A =-3?

OpenStudy (lgbasallote):

yeah that's right

OpenStudy (anonymous):

ok cool i was about to write out how i got it lol

OpenStudy (lgbasallote):

everything's right

OpenStudy (anonymous):

and now i just plug back all the variables right

hartnn (hartnn):

yup, and use inverse laplace formulas.

OpenStudy (lgbasallote):

yep find \[\huge \mathcal L ^{-1} \left \{ -\frac 3s + \frac 5s + \frac 4{s+1} \right \}\]

OpenStudy (lgbasallote):

uhh that should be 5/s^2

OpenStudy (anonymous):

ok sorry im just learning this but how do you go about doing that im trying to find the L^-1 formula but my notes arent really helping

OpenStudy (lgbasallote):

here's a hint: \[\huge \mathcal L \{ a \} = \frac as\] \[\huge \mathcal L \{ t^n\} = \frac{n!}{s^{n+1}}\] \[\huge \mathcal L \{e^{at} \} = \frac{a}{s -a}\] does that help?

OpenStudy (lgbasallote):

uhhh wait... wrong formula for last one

OpenStudy (lgbasallote):

\[\huge \mathcal L \{e^{at} \} = \frac{1}{s-a}\] better

OpenStudy (anonymous):

yes thats very helpful but im getting confused with what n should be in the middle equations

OpenStudy (lgbasallote):

n is the exponent

OpenStudy (anonymous):

so it would be 2?

OpenStudy (lgbasallote):

for example: \[\huge \mathcal L \{ t^2 \} = \frac{2!}{s^{2+1}} \implies \frac{2}{s^3}\]

OpenStudy (anonymous):

thats whats a lil confusing to me because on the denominator it would have to be n=1 so that it would equal s^2 but on the numerator 1! equals 1 right ?

hartnn (hartnn):

yes thats true, so u will have 5t^1= 5t ....isn't it ? u can verify that u should get final answer as 5t-3+4e^(-t) did u get that?

OpenStudy (anonymous):

yup thats what i got thanks

hartnn (hartnn):

welcome :)